Compare simple and compound interest (compounded annually) on a principal of $25,000 at an annual rate of 4%. Find the number of years it takes for the the difference of the TOTAL using Simple interest and TOTAL using compound interest (compounded annually) to reach more than $1000.
Answer #1

Solution :-

As we know the Simple interest is charged always on the Principal Amount so the Simple interest rate per year is Same

Therefore Simple interest every year = $25,000 * 4% = $1,000

And in Case of Compounding interest the interest rate is charged on the outstanding balance that is Principal Amount and interest due there on.

Compound Interest in first Year = $25000 * 4% = $1000

Balance due after 1st year = $25000 + $1000 = $26000

Compound Interest in Second Year = ( $25000 + $1000 ) * 4% = $1040

Balance due after second year = $26000 + $1040 = $27040

And So on.....

C D Τ Ε F H36 B 1 Compound Interest Table 2 Year Opening Balance 3 1 25000 2 26000 3 27040 28121.60 29246.46 30416.32 31632.9

- fx H36 А 1 Compound 2 Year E 31 Opening Balance -25000 =D3 =D4 -D5 OU Interest =B3*4% =B4*4% =B5*4% =B6*4% =B7*4% =B8*4% =B

the number of years it takes for the the difference of the TOTAL using Simple interest and TOTAL using compound interest (compounded annually) to reach more than $1000 is 8 Years

If there is any doubt please ask in comments